OpenStudy (anonymous):

hi guys , can any one help me with this ? Atruck accelerates from rest at 0.3m/s for 30 sec,continues at constant speed for 2 min , and then comes to a stop in 15 sec , What distance did it cover ?

OpenStudy (anonymous):

in first 30 sec,S1=135m in next 2 min,S2=1080m in final 15 sec,S3=67.5m total distance covered,s=s1+s2+s3=1282.5

OpenStudy (anonymous):

Thanks , what law did u use to get s1 , s2 ,s3 ?

OpenStudy (anonymous):

for s1 and s2,i used the equation:s=ut+(at.t)/2 for s3 i used : s=(average velocity)xtime